OpenStudy (anonymous):

If f(x) = 7x – 12, what is f–1(x) ?

OpenStudy (bahrom7893):

y = 7x - 12

OpenStudy (bahrom7893):

solve for y: y + 12 = 7x x = (y+12)/7

OpenStudy (bahrom7893):

so now just swap x and y: y = (x+12)/7

OpenStudy (bahrom7893):

so f-1(x) = (x+12)/7

OpenStudy (anonymous):

F(x)=7x-12 Y=7x-12 X+12=7y x+12/7=y y+12/7=x=f-1(x)